This is a differential equations course, please answer all the questions.

Find the general solution.

1 - 1 8. I(t) = 3 3 2 2 1 -1 (t) 3 2 -2 9. z(t) -2)=(1) 10. 2(7) = (1 -1)=(0) 11. z(t) 2(t) =) =(0) 13. zº(t) = 6 (5 -3)
Answer #1

4. (t) x (1) 2 2 1 4 ARCL) Where A = 3 2. 2 1 I- -1 4. IA - 5) 3 2.) al 2 T =0 (TX) raft (1-1) [41-4) (2-2) + 1] + [41-1)3-(A²-7-1) + 5 72-a1-5 =0 -) 1?-1-6 = 0 > 32-34 of 21-6 y yo t (4-3) 42 (0-3) = 0 - x = 3,12 a = 3 1 2 =1 3 = -2 are Valuel TTfor X 3. x (A -x) x 1-d 4 2-) k3 2 4 4 T 3. 2-3 2 23 2 1 1 -1-3 4 4 3 22 -1 1 13 2 1 - 4 R > 2 R3+3R2 R3 → R3+ R2 -14 0 -5 x2let x3 =1 V 2 for Xal (A-2)* :o 1-) 4 3 2- 12 23 o 2 1 -1-) T 94 3 2 - - K2 co 2 23 1 -1 4 3 000 22 no 2 . 2 .2 3 1 U2 O 0 4- 12t 423 0 커 기2 403 2죄 + K2 - 2X3 이 - 2의+ 423 ~2X y 24 4243 = 3 23 1 20 43 23 4 23 23 ht 3. 디 V 4 for A2 (A -Ali) 4 ㅠ 3 2 어2R2 R2 R R3 - 3R2 - 281 3 4. 5 -5 ( 23 -5 3 4 M M 0 5 -5 X2 43 5 H - 583 l2 =M3. 34 - x2 + 443 34 - k3 + 483 - 3er +343=0 = -2Soln of 0 is nivou by. x (t) = arietet quiet & cauzet + C3 V3 e at + C2 Vz edzt & Ca va etzt X(t) G 3t e & C2 2 + at e + whic